OpenStudy (anonymous):

What is the equation of the function y=2/x translated 3 units to the left and 4 units up?

OpenStudy (mathmale):

Translating the graph of y=2/x four units up results in a new function: h(x)=2/x + 4. Have you tried looking up "horizontal translation?"

OpenStudy (anonymous):

If its translated to the left would be y=2/x-3+4 ? Am i right ?

OpenStudy (anonymous):

okay thank you so much

OpenStudy (mathmale):

@csmath: great that you've pointed out that parentheses are missing! Please, would you write out the rules for horizontal translation, and focus especially on which sign to use: + or -.

OpenStudy (mathmale):

Driana, please work with csmath towards agreement on what the final answer should be.

OpenStudy (anonymous):

ohh okay that makes sense, thank you <3